Tom's father is 3 times as old as Tom. 4 years ago, he was 4 times older. How old is Tom?
Answer by chitra(359)

Let the age of Tom be "x" yrs.

Therefore, the age of Tom's father is 3x yrs.

Given that 4 yrs ago, he was 4 times as old as tom. Therfore the equations are:

3x - 4 = 4(x - 4)

3x - 4 = 4x - 16

16 - 4 = 4x - 3x

12 = x

Therfore, Tom is 12 yrs old

Hence his father would be 3 times as tom. that is equal to 3(12) = 36 yrs

Hence the solution

regs